OpenStudy (winsteria):

Jane walked 4/5 mile on Monday. On Tuesday, she walked 1/2 of the distance she walked on Monday. How many miles did Jane walk on Tuesday? A) 3/4 mile B) 1/2 mile C) 1/5 mile D) 2/5 mile

OpenStudy (winsteria):

@KamiBug

OpenStudy (winsteria):

I think it's C ._.

OpenStudy (kamibug):

4/5 * 1/2 = 4/10 = 2/5

OpenStudy (anonymous):

she walked a quarter of the distance she walked on monday.... So she has walked 2 times less....

OpenStudy (winsteria):

Thank you both

OpenStudy (anonymous):

NP